Grade 12th passAlgebra

Four dice are rolled . The number of ways in which at least one die shows 3 is

Akansha yadav
8 Years agoGrade 12th pass

2 Answers

Anish Singhal
7 Years ago

No. of ways in which any number appearing in one dice = 6
No. of ways in which 3 appear in one dice = 1
No. of ways in which 3 does not appear in one dice = 5
There are 4 dice.
Getting 3 in atleast one dice =
Getting any number in all the 4 dice - Getting not 3 in any of the 4 dice.
=(6×6×6×6)−(5×5××5×5)=(6×6×6×6)−(5×5××5×5)
1296−625=671
Anish Singhal
7 Years ago

No. of ways in which any number appearing in one dice = 6
No. of ways in which 3 appear in one dice = 1
No. of ways in which 3 does not appear in one dice = 5
There are 4 dice.
Getting 3 in atleast one dice =
Getting any number in all the 4 dice - Getting not 3 in any of the 4 dice.
=(6×6×6×6)−(5×5××5×5)=(6×6×6×6)−(5×5××5×5)
1296−625=671